OpenStudy (meilendurcer):

how to solve 4/n+2 = 7/n

OpenStudy (nikato):

Cross multiply. a/b = c/d Becomes a x d= b x c

OpenStudy (nikato):

So 4 x n = (n+2) 7

OpenStudy (nikato):

Distribute the right side (n+2)7 =n(7) + 2(7) = 7n+14

OpenStudy (nikato):

So 4n= 7n+14 And try to solve for n from here...

OpenStudy (meilendurcer):

i can't find n

OpenStudy (nikato):

-3n=14 Divide both sides by -3 n=14/-3